What is the maximum amount in moles of P2O5 that can theoretically be made from 112 g of P4 and excess oxygen?
1
Expert's answer
2019-01-29T05:02:26-0500

P4 + 5O2 = 2P2O5

n= m/M

M(P4)= 30.97*4 = 123.88 g/mol

n(P4)= 112g/123.88 g/mol = 0.904 mol

According to equation 1 mole of P4 gives 2 moles of P2O5

We have 0.904 mol of P4 which give x mol of P2O5

Solve the proportion:

1/0.904 = 2/x

x= 1.808

n(P2O5) = 1.808 mol

Answer: 1.808 mol